OpenStudy (anonymous):

A 16-inch candle is lit and burns at a constant rate. After 30 minutes, the candle is 12 inches tall. Which linear equation models the relationship between h, the height of the candle in inches, and t, the time in hours that it has burned? a. 8t + h = 16 b. 8t – h = 16 c. 4t + 3h = 16 d. 4t – 3h = 16

OpenStudy (anonymous):

@Loser66

OpenStudy (texaschic101):

I am trying to wake up here..so please be patient..lol. The candle starts at 16 in....after 30 min the candle is 12 in....this means that the candle lost 4 inches for every 30 min, or 8 inches for every hr. h (height) ------ t (hours) 12 1/2 8 1 4 1.5 0 2 now lets find your equation.. 8t + h = 16 8(1) + 8 = 16 8 + 8 = 16 16 = 16 I am going with 8t + h = 16...but you might want to get a second opinion..I am not completely sure about this.

OpenStudy (anonymous):

@bibby Is she Correct ?

OpenStudy (texaschic101):

@primeralph ...help